At STP a certain mass of gas occupies a volume of 800cm3.find the temperature at which the gas occupies 940cm3 and has a pressure of mmttg


1
Expert's answer
2021-02-01T03:59:28-0500

Standard Temperature and Pressure (STP) is defined as 0 degrees Celsius and 1 atmosphere of pressure.

"T_1 = 0 = 273 \;K \\

V_1 = 800 \;cm^3 \\

P_1 = 1 \;atm \\

V_2 = 940 \;cm^3"

Let’s use the value of "P_2 = 1520 \;mmHg = 2 \;atm \\"

"\frac{P_1 \times V_1}{T_1} = \frac{P_2 \times V_2}{T_2} \\

T_2 = \frac{T_1 \times P_2 \times V_2}{P_1 \times V_1} \\

= \frac{273 \times 2 \times 940}{1 \times 800} \\

= 641.55 \;K \\

T_2 = 368 \;ºC"